# for the Hall C SIDIS Collaboration

### v1

@@ Abstract
Measurements of SIDIS multiplicities for $\pi^{+}$ and $\pi^{-}$ from proton and deuteron targets are reported on a grid of hadron kinematic variables $z$, $P_{T}$, and $\phi^{*}$ for leptonic kinematic variables in the range $0.3<x<0.6$ and $3<Q^{2}<5$ GeV2. Data were acquired in 2018-2019 at Jefferson Lab Hall C with a 10.6 GeV electron beam impinging on 10-cm-long liquid hydrogen and deuterium targets. Scattered electrons and charged pions were detected in the HMS and SHMS spectrometers, respectively. The multiplicities were fitted for each bin in $(x,~Q^{2},~z,~P_{t})$ to extract the $\phi^{*}$ independent $M_{0}$ and the azimuthal modulations $\langle\cos(\phi^{*})\rangle$ and $\langle\cos(2\phi^{*})\rangle$. The $P_{t}$-dependence of the $M_{0}$ results was found to be remarkably consistent for the four cases studied: $ep\rightarrow e\pi^{+}X$, $ep\rightarrow e\pi^{-}X$, $ed\rightarrow e\pi^{+}X$, $ed\rightarrow e\pi^{-}X$ over the range $0<P_{t}<0.4$ GeV, as were the multiplicities evaluated near $\phi^{*}=180^{\circ}$ over the extended range $0<P_{t}<0.7$ GeV. The Gaussian widths of the $P_{t}$-dependence exhibit a quadratic increase with $z$. The $\cos(\phi^{*})$ modulations were found to be consistent with zero for $\pi^{+}$, in agreement with previous world data, while the $\pi^{-}$ moments were, in many cases, significantly greater than zero. The $\cos(2\phi^{*})$ modulations were found to be consistent with zero. The higher statistical precision of this dataset compared to previously published data should allow improved determinations of quark transverse momentum distributions and higher twist contributions.
@@ I Introduction
Over the last five decades, semi-inclusive deep-inelastic (SIDIS) lepton-nucleon scattering ($lN\rightarrow l^{{}^{\prime}}hX$) has proven to be a key tool in building a more complete and accurate picture of the internal structure of the nucleon in terms of the partonic degrees of freedom of quantum chromodynamics (QCD). It has been instrumental in establishing that the collinear picture of the quark-parton model is incomplete. One of the most important advantages of SIDIS is the ability to measure the yield of hadrons ($h$) both in terms of the longitudinal momentum fraction $z$ and the transverse momentum $P_{t}$ (shown schematically in Fig. 1). The SIDIS process in its simplest interpretation can be thought of as a subset of deep-inelastic scattering (DIS), described by parton distribution functions (PDFs), with a multiplicity function ($M$) that indicates the probability of the DIS final-state containing a particular meson with a particular momentum vector. In this highly simplified picture, the multiplicity dependence on $P_{t}$ arises from a convolution of the transverse momentum of the quark ($k_{T}$) and the transverse momentum generated in the fragmentation process ($p_{\perp}$), in which the struck quark hadronizes into multiple final-state particles. A comparison of SIDIS from protons ($u$-quark dominated) and neutrons ($d$-quark dominated) could, in principle, be used to constrain the difference between the average $k_{T}$ of up and down valence quarks in the nucleon. Expanding the kinematic coverage for both positive and negative pions can help to distinguish differences in “favored” and “unfavored” fragmentation functions, where “favored” refers to a pion containing the struck quark. The azimuthal modulations of the measured pion relative to the virtual photon direction are also sensitive to $k_{T}$, especially when the incident electron or the target nucleon are polarized. Anselmino et al. (2005, 2006)
@@ I.1 Formalism
The semi-inclusive scattering of unpolarized electrons by unpolarized nucleons in the SIDIS kinematic region can be described formally Bacchetta et al. (2007) in terms of structure functions (SF) as
@@ I.2 Theoretical interpretation
Although the intrinsic transverse momentum of the partons, $\vec{k}_{T}$, was already considered by Feynman when he introduced the parton model Feynman (1969), the initial experimental and theoretical focus was on the one-dimensional light-cone momentum variable $x$. In recent years significant advances have been made in incorporating $\vec{k}_{T}$ into the theoretical description of SIDIS processes. For example, the transverse momentum dependent (TMD) parton distribution functions (PDF) and fragmentation functions (FF) Mulders and Tangerman (1996); Boer et al. (1998) were introduced, and a TMD factorization formalism Ji et al. (2004) was developed. The factorization framework demonstrates that the hadron transverse momentum arises from the transverse momentum of the quarks in the nucleon, convoluted with the transverse momentum generated during quark fragmentation. Both the TMD PDF and the FF depend on two independent variables: the TMD on $x$ and $k_{T}$, while the FF depends on $z$ and the transverse momentum $p_{\perp}$ of the hadron acquired during the fragmentation process. The TMD factorization was first shown for the high-energy limit (high values of the virtuality scale, $Q^{2}>>\Lambda_{\text{QCD}}$) and moderate values of $P_{t}\sim\Lambda_{\text{QCD}}$. However, its applicability at moderate $Q^{2}$ (2 - 4 GeV2) has since been observed in several experiments Navasardyan et al. (2007); Mkrtchyan et al. (2008); Osikenki et al. (2009); Asaturyan et al. (2012). Within this framework, and with the approximation that higher-order (higher-twist) corrections are suppressed by powers of $1/Q$, the SIDIS differential cross section for polarized leptons on polarized nucleons is expressed in terms of 18 structure functions that are convolutions of various TMD PDF and FF Bacchetta et al. (2007). This large number of structure functions is a consequence of the fact that, for a spin-1/2 hadron, there are 8 TMD Mulders and Tangerman (1996); Bacchetta et al. (2007); Anselmino et al. (2011), each representing a unique correlation between the spin and the orbital motion of the partons. These TMD are parameterized using the world data on SIDIS and other processes Barone et al. (2010); Bacchetta and Radici (2011); Anselmino et al. (2014); Barone et al. (2015); Bacchetta et al. (2017).
@@
As expected, the unpolarized SIDIS cross section can only provide information about the unpolarized TMD distribution functions and the unpolarized TMD fragmentation functions. The $\cos(\phi^{*})$ dependence was predicted in 1978 by R. Cahn Cahn (1978) as a result of the interaction of the virtual photon with quarks in the nucleon possessing intrinsic transverse momentum. Both the $\cos(\phi^{*})$ and $\cos(2\phi^{*})$ modulations receive contributions from the Boer-Mulders effect Boer and Mulders (1998), arising from a correlation between the quark’s intrinsic transverse momentum and its transverse spin, coupled to the Collins fragmentation function Collins (1993), which preserves the correlation with fragmentation dependent on the struck quark’s transverse spin. Phenomenological analyses by Barone et al. Barone et al. (2015) stress that these structure functions are sensitive to higher-twist contributions. Additionally, the transverse momentum dependence of the TMD and FF are expected to be approximately Gaussian Anselmino et al. (2005), for low values of $P_{t}$. To leading order, this simplification and momentum conservation give: $\langle\vec{P}^{2}_{t}\rangle\simeq\langle\vec{p}^{2}_{\perp}\rangle+z^{2}\langle\vec{k}^{2}_{T}\rangle$, implying that the transverse momentum dependence of TMD and FF can be parameterized by a normalized linear combination of a Gaussian and a $z^{2}$-weighted Gaussian Bacchetta et al. (2017). This interpretation relies on $\langle\vec{p}^{2}_{\perp}\rangle$ being independent of $z$, which is manifestly incorrect as $z\rightarrow 0$ at finite values of $\nu$.
@@ I.3 Previous experiments
Some of the earliest SIDIS experimental studies in the valence quark region ($x>0.25$) were made at Cornell in the 1970s, using 12 GeV electrons Bebek et al. (1975). These experiments demonstrated that multiplicities behave roughly as $(1-z)^{2}$ for $z<0.7$, have an approximately Gaussian distribution in $P_{t}$, and have relatively small dependence on $\phi^{*}$ compared to exclusive pion electroproduction. Subsequent experiments Ashman et al. (1991); Navasardyan et al. (2007); Osikenki et al. (2009); Airapetian et al. (2013a); Adolph et al. showed that spin-averaged cross sections can be described as a convolution of quark PDFs derived from DIS and Drell-Yan reactions with FF derived from electron-positron colliders. They also demonstrated the usefulness of describing the production of the leading meson that contains the struck quark flavor with “favored” FF, while other mesons are described by “unfavored” FF, which exhibit a smaller strength at high $z$ than favored FF. An experiment at JLab Hall C with 6 GeV electrons Navasardyan et al. (2007) showed a duality between the results in the resonance region, at low center-of-mass energy ($2<W<3$ GeV), and the results well above the nucleon resonance region ($W>3$ GeV). This was true as long as the electron-pion invariant missing mass squared, $M_{x}^{2}$, was well above 2.5 GeV2 (corresponding to $z<0.7$ at these kinematic settings), as shown in Fig. 2. The noticeable peak centered at $M_{x}^{2}=1.5$ GeV2, visible in this figure, was due to the semi-exclusive channel $ep\rightarrow e\pi\Delta(1232)$, which was not subtracted in that analysis (but is subtracted in the present analysis). Simple phenomenological fits Mkrtchyan et al. (2008); Asaturyan et al. (2012) to these data attempted to disentangle the up and down valence quark $k_{T}$ widths, as well as favored and unfavored FF widths, with the assumptions that the $\cos(\phi^{*})$ dependence is dominated by the Cahn term and that the fragmentation widths are independent of $z$ (both of which have since been shown to be incorrect).
Figure 2: SIDIS cross sections for charged pions from proton and deuteron targets compared to Monte Carlo simulations (dashed lines) from Ref. Navasardyan et al. (2007) with 5.6 GeV electrons at JLab Hall C as a function of $z$ and corresponding electron-pion missing mass squared $M_{x}^{2}$.
The experimental results prior to 2018 cannot be considered conclusive due to the limited kinematic coverage, low counting rates, inadequate particle identification, and poor resolution in $\phi^{*}$ at low $P_{T}$. In order to overcome many of these limitations, a new experimental program was initiated at the energy-upgraded JLab, using both the wide-acceptance, lower luminosity CLAS12 detector in Hall B and the high-luminosity, small acceptance spectrometers in Hall C. The broad program includes the use of beam and target polarization, both light and heavy nuclear targets, a range of electron beam energies, and identification of many final-state mesons. In this paper, we report on spectrometer-based results for charged pions with an unpolarized beam and target at the highest available beam energy, from an experiment Mkrtchyan et al. that was an integral part of the JLab SIDIS program and was completed in 2019. The experiment featured a wide range of $(x,Q^{2})$ values (to study higher-twist contributions), full $\phi^{*}$ coverage for $P_{T}<0.25$ GeV, a larger $P_{T}$ range for $\phi^{*}$ near 180, and a broad range in $z$ (to help distinguish $k_{T}$ width from $p_{\perp}$ widths).
@@ II The Experiment
The experiment was carried out in spring 2018 and fall 2019, in Hall C at JLab. Electrons scattered from hydrogen and deuterium cryogenic targets were detected in the High Momentum Spectrometer (HMS), with alternatively positive and negative pions detected in the Super High Momentum Spectrometer (SHMS). An overview of the experiment layout is shown in Fig. 3 S. Ali et al. .
@@ II.4 Electron identification
Scattered electrons were detected in the well-studied High Momentum Spectrometer SURA-CEBAF (1990), which has been in use since 1996. As shown in Fig. 4, the detector package includes two pairs of segmented planes of plastic scintillators used to determine the time of arrival of particles transported through the spectrometer magnets, with a timing resolution of about 0.3 ns. The planes were grouped into pairs (x, y) separated by 2 m, allowing the particle speed to be determined with a resolution of about 0.1 in $\beta$.
Figure 4: Detector configuration in the HMS
@@
A two-mirror threshold gas Cherenkov detector and a segmented Pb-glass calorimeter Mkrtchyan et al. (2013) were used to distinguish electrons from pions (both of which generally passed the cut on scintillator paddles). The Cherenkov detector gas mixture and pressure were set to give a pion momentum threshold of 4.5 GeV. The average number of photo-electrons (p.e.) produced by electrons was about 6 (10) in the upper (lower) mirrors for settings I and II, due to a crack in the upper mirror. With a threshold of 0.3 p.e., detection efficiency was determined to be $99.5\%$. For setting III, the cracked mirror was fixed, resulting in an average of 10 photoelectrons from each mirror. A light leak resulted in very high counting rates, which effectively blocked some of the electron signals from being read out. Using a threshold of 1 p.e. to reduce this effect, we determined an effective average efficiency of 0.975 for the electrons, independent of beam current.
Scattered electrons were identified in the segmented lead glass electromagnetic calorimeter using the ratio of energy deposited in the blocks near the projected track position ($E_{cal}$) to the track momentum ($P_{e}$). As illustrated in Fig. 5, the $E_{cal}/P_{e}$ distributions for each of the three kinematic settings show a narrow peak centered on unity. The small flat-looking distributions below 0.75 are mostly from pions. The vertical dashed line at 0.75 shows the cut used for electron identification. The electron detection efficiency of the HMS calorimeter was greater than 99.7%. After correcting for accidental coincidences, the contamination of pions in the final event sample was less than 0.5%.

### v2

@@ Abstract
Measurements of SIDIS multiplicities for $\pi^{+}$ and $\pi^{-}$ from proton and deuteron targets are reported on a grid of hadron kinematic variables $z$, $P_{T}$, and $\phi^{*}$ for leptonic kinematic variables in the range $0.3<x<0.6$ and $3<Q^{2}<5$ GeV2. Data were acquired in 2018-2019 at Jefferson Lab Hall C with a 10.6 GeV electron beam impinging on 10-cm-long liquid hydrogen and deuterium targets. Scattered electrons and charged pions were detected in the HMS and SHMS spectrometers, respectively. The multiplicities were fitted for each bin in $(x,~Q^{2},~z,~P_{t})$ to extract the $\phi^{*}$ - independent $M_{0}$ and the azimuthal modulations $\langle\cos(\phi^{*})\rangle$ and $\langle\cos(2\phi^{*})\rangle$. The $P_{t}$-dependence of the $M_{0}$ results was found to be remarkably consistent for the four cases studied: $ep\rightarrow e\pi^{+}X$, $ep\rightarrow e\pi^{-}X$, $ed\rightarrow e\pi^{+}X$, $ed\rightarrow e\pi^{-}X$ over the range $0<P_{t}<0.4$ GeV, as were the multiplicities evaluated near $\phi^{*}=180^{\circ}$ over the extended range $0<P_{t}<0.7$ GeV. The Gaussian widths of the $P_{t}$-dependence exhibit a quadratic increase with $z$. The $\cos(\phi^{*})$ modulations were found to be consistent with zero for $\pi^{+}$, in agreement with previous world data, while the $\pi^{-}$ moments were, in many cases, significantly greater than zero. The $\cos(2\phi^{*})$ modulations were found to be consistent with zero. The higher statistical precision of this dataset of about 20,000 individual multiplicity values, compared to previously published data, should allow improved determinations of quark transverse momentum distributions and higher twist contributions.
@@ I Introduction
Over the last five decades, semi-inclusive deep-inelastic (SIDIS) lepton-nucleon scattering ($lN\rightarrow l^{{}^{\prime}}hX$) has proven to be a key tool in building a more complete and accurate picture of the internal structure of the nucleon in terms of the partonic degrees of freedom of quantum chromodynamics (QCD). It has been instrumental in establishing that the collinear picture of the quark-parton model is incomplete. One of the most important advantages of SIDIS is the ability to measure the yield of hadrons ($h$) both in terms of the longitudinal momentum fraction $z$ and the transverse momentum $P_{t}$ (shown schematically in Fig. 1). The SIDIS process in its simplest interpretation can be thought of as a subset of deep-inelastic scattering (DIS), described by parton distribution functions (PDFs), with a multiplicity function ($M$) that indicates the probability of the DIS final-state containing a particular meson with a particular momentum vector. In this highly simplified picture, the multiplicity dependence on $P_{t}$ arises from a convolution of the transverse momentum of the quark ($k_{T}$) and the transverse momentum generated in the fragmentation process ($p_{\perp}$), in which the struck quark hadronizes into multiple final-state particles. A comparison of SIDIS from protons ($u$-quark dominated) and neutrons ($d$-quark dominated) could, in principle, be used to constrain the difference between the average $k_{T}$ of up and down valence quarks in the nucleon. Expanding the kinematic coverage for both positive and negative pions can help to distinguish differences in “favored” and “unfavored” fragmentation functions, where “favored” refers to a pion containing the struck quark. The azimuthal modulations of the measured pion relative to the virtual photon direction are also sensitive to $k_{T}$, especially when the incident electron or the target nucleon are polarized. ansel05; ansel06
@@ I.1 Formalism
The semi-inclusive scattering of unpolarized electrons by unpolarized nucleons in the SIDIS kinematic region can be described formally Bacchetta07 in terms of structure functions (SF) as
@@ I.2 Theoretical interpretation
Although the intrinsic transverse momentum of the partons, $\vec{k}_{T}$, was already considered by Feynman when he introduced the parton model FF1, the initial experimental and theoretical focus was on the one-dimensional light-cone momentum variable $x$. In recent years significant advances have been made in incorporating $\vec{k}_{T}$ into the theoretical description of SIDIS processes. For example, the transverse momentum dependent (TMD) parton distribution functions (PDF) and fragmentation functions (FF) Muld96; Boer98 were introduced, and a TMD factorization formalism Ji2004 was developed. The factorization framework demonstrates that the hadron transverse momentum arises from the transverse momentum of the quarks in the nucleon, convoluted with the transverse momentum generated during quark fragmentation. Both the TMD PDF and the FF depend on two independent variables: the TMD on $x$ and $k_{T}$, while the FF depends on $z$ and the transverse momentum $p_{\perp}$ of the hadron acquired during the fragmentation process. The TMD factorization was first shown for the high-energy limit (high values of the virtuality scale, $Q^{2}>>\Lambda_{\text{QCD}}$) and moderate values of $P_{t}\sim\Lambda_{\text{QCD}}$. However, its applicability at moderate $Q^{2}$ (2 - 4 GeV2) has since been observed in several experiments Tigran07; Hamlet08; Osipenko09; Asatur12. Within this framework, and with the approximation that higher-order (higher-twist) corrections are suppressed by powers of $1/Q$, the SIDIS differential cross section for polarized leptons on polarized nucleons is expressed in terms of 18 structure functions that are convolutions of various TMD PDF and FF Bacchetta07. This large number of structure functions is a consequence of the fact that, for a spin-1/2 hadron, there are 8 TMD Muld96; Bacchetta07; Ancel11, each representing a unique correlation between the spin and the orbital motion of the partons. These TMD are parameterized using the world data on SIDIS and other processes Barone10; Bacchetta11; Ancel14; Barone:2015ksa; Bacchetta17.
@@
As expected, the unpolarized SIDIS cross section can only provide information about the unpolarized TMD distribution functions and the unpolarized TMD fragmentation functions. The $\cos(\phi^{*})$ dependence was predicted in 1978 by R. Cahn Cahn78 as a result of the interaction of the virtual photon with quarks in the nucleon possessing intrinsic transverse momentum. Both the $\cos(\phi^{*})$ and $\cos(2\phi^{*})$ modulations receive contributions from the Boer-Mulders effect BoerMuldersTimeR, arising from a correlation between the quark’s intrinsic transverse momentum and its transverse spin, coupled to the Collins fragmentation function COLLINS1993161, which preserves the correlation with fragmentation dependent on the struck quark’s transverse spin. Phenomenological analyses by Barone et al. Barone:2015ksa stress that these structure functions are sensitive to higher-twist contributions. Additionally, the transverse momentum dependence of the TMD and FF are expected to be approximately Gaussian ansel05, for low values of $P_{t}$. To leading order, this simplification and momentum conservation give: $\langle\vec{P}^{2}_{t}\rangle\simeq\langle\vec{p}^{2}_{\perp}\rangle+z^{2}\langle\vec{k}^{2}_{T}\rangle$, implying that the transverse momentum dependence of TMD and FF can be parameterized by a normalized linear combination of a Gaussian and a $z^{2}$-weighted Gaussian Bacchetta17. This interpretation relies on $\langle\vec{p}^{2}_{\perp}\rangle$ being independent of $z$, which is manifestly incorrect as $z\rightarrow 0$ at finite values of $\nu$.
@@ I.3 Previous experiments
Some of the earliest SIDIS experimental studies in the valence quark region ($x>0.25$) were made at Cornell in the 1970s, using 12 GeV electrons Cornell, with an integrated luminosity several orders of magnitude lower than the present experiment with 10.6 GeV electrons. These experiments demonstrated that multiplicities behave roughly as $(1-z)^{2}$ for $z<0.7$, have an approximately Gaussian distribution in $P_{t}$, and have relatively small dependence on $\phi^{*}$ compared to exclusive pion electroproduction. Subsequent experiments by the HERMES collaboration HERMES:2012uyd; HERMES:2012kpt used 27 GeV electrons or positrons scattering from very thin targets at DESY. The results are displayed as a function of $z$ averaged over $(x,Q^{2},P_{t})$ and a function of $P_{t}$ averaged over $(x,Q^{2},z)$, although a later publication Barone:2015ksa displays azimuthal asymmetries averaged over $z$ in some bins in $(x,Q^{2})$. The HERMES results are all at a bit higher values of $W$ than the present experiment.
Experiments at CERN by the EMC EMC_PT_ref and later by the COMPASS collaboration COMPASS:2024gje; COMPASS_PT_ref collaboration used 100-200 GeV muon beams and focused on the sea quark region ($x<0.3$). The most recent COMPASS publications show multiplicities as a function of $z$ averaged over $P_{t}$ COMPASS:2024gje, and azimuthal dependencies as a function of $P_{t}$ averaged over $x$ and $z$ Benesova:2024cfk. Earlier publications COMPASS_PT_ref are for un-separated hadrons from a nuclear target (LiD). All of the CERN results are at considerably higher values of $W$ than the present experiment.
Global fits to the HERMES and COMPASS multiplicity results have shown that spin-averaged cross sections can be reasonably well described as a convolution of quark PDFs derived from DIS and Drell-Yan reactions with fragmentation functions (FF) derived from electron-positron colliders. They also demonstrated the usefulness of describing the production of the leading meson that contains the struck quark flavor with “favored” FF, while other mesons are described by “unfavored” FF, which exhibit a smaller strength at high $z$ than favored FF. In this paper, we compare our multiplicity results to a very recent global fits to the HERMES and COMPASS multiplicity results by the MAP collaboration MAP24, and to the the azimuthal asymmetries by Barone et al. Barone:2015ksa.
[…]
@@ II The Experiment
The experimental results prior to 2018 cannot be considered conclusive in the quark valence region ($x>0.25$) due to the limited kinematic coverage, low counting rates, inadequate particle identification, and poor resolution in $\phi^{*}$ at low $P_{T}$. In order to overcome many of these limitations, a new experimental program was initiated in 2018 using beam energies up to 11 GeV with both the wide-acceptance, lower luminosity CLAS12 detector in Hall B and the high-luminosity, small acceptance spectrometers in Hall C. The broad program includes the use of beam and target polarization, both light and heavy nuclear targets, a range of electron beam energies, and identification of many final-state mesons. In this paper, we report on spectrometer-based results for charged pions with an unpolarized target at the highest available beam energy, from an experiment that was an integral part of the JLab SIDIS program and was completed in 2019. The experiment featured a wide range of $(x,Q^{2})$ values (to study higher-twist contributions), full $\phi^{*}$ coverage for $P_{t}<0.3$ GeV, a larger $P_{t}$ range for $\phi^{*}$ near 180, and a broad range in $z$ (to help distinguish $k_{T}$ width from $p_{\perp}$ widths). The multiplicity results were published as a function of $W$ integrated over $\phi^{*}$ and $0<P_{t}<0.2$ GeV in a recent Letter Bhatt:2024prq. In this paper, we present results in a fully 3-dinemsional grid in $(z,P_{t},\phi^{*}$ for three values of $(x,Q^{2}$, spanning $2.6<W<3.3$ GeV.
The experiment was carried out in spring 2018 and fall 2019, in Hall C at JLab. Electrons scattered from hydrogen and deuterium cryogenic targets were detected in the High Momentum Spectrometer (HMS), with alternatively positive and negative pions detected in the Super High Momentum Spectrometer (SHMS). Additional details about the experiment can be found in Refs. Bhatt:2024prq; shms_nim. An overview of the experiment layout is shown in Fig. 2.
@@ II.4 Electron identification
Scattered electrons were detected in the well-studied High Momentum Spectrometer hms_nim, which has been in use since 1996.
@@
A two-mirror threshold gas Cherenkov detector and a segmented Pb-glass calorimeter cal_ref were used to distinguish electrons from pions (both of which generally passed the cut on scintillator paddles). The Cherenkov detector gas mixture and pressure were set to give a pion momentum threshold of 4.5 GeV. Detection efficiency was determined to be $99.5\%$ for settings I and II. For setting III, the average efficiency was 97.5% for the electrons, independent of beam current.
Scattered electrons were identified in the segmented lead glass electromagnetic calorimeter using the ratio of energy deposited in the blocks near the projected track position ($E_{cal}$) to the track momentum ($P_{e}$). As illustrated in Fig. 3, the $E_{cal}/P_{e}$ distributions for each of the three kinematic settings show a narrow peak centered on unity. The vertical dashed line at 0.75 shows the cut used for electron identification. The electron detection efficiency of the HMS calorimeter was greater than 99.7%. After correcting for accidental coincidences, the contamination of pions in the final event sample was less than 0.5%.